Take all possible samples of size 3 without replacement from population 0, 3, 6, 8, 9, 12, 15. Make a sampling distribution of Proportion odd numbers. Find its mean and Standard Deviation. Verify the properties.


1
Expert's answer
2021-08-09T13:07:58-0400

N= 7

n=3

Number of samples "= \frac{N!}{n!(N-n)!}"

"= \frac{7!}{3!(7-3)!} = \frac{210}{6}=35"



The sampling distribution of Proportion odd numbers


Mean x "= \sum x \times P(x) = 0.428"

"SD = \sqrt{\sum (x^2 \times P(x) -(\sum x \times P(x))^2} \\

= \sqrt{0.238 -(0.428)^2} \\

= \sqrt{0.238-0.183} \\

= \sqrt{0.055} \\

=0.234"